sxdsouls:

help besties Jacob’s age is two years more than the sum of the ages of his siblings Becky and Micah. Which equation represents Jacob’s age?

sxdsouls:

A. z = x + y − 2; x represents Micah's age, y represents Becky's age, and z represents Jacob's age B. x = y + z + 2; x represents Jacob's age, y represents Micah's age, and z represents Becky's age C. x = 2 + y + z; x represents Becky's age, y represents Jacob's age, and z represents Micah's age D. y = x + z − 2; x represents Jacob's age, y represents Becky's age, and z represents Micah's age

sxdsouls:

nevermind b byeeeeeeeeeeefcr

AZ:

@sxdsouls wrote:
nevermind b byeeeeeeeeeeefcr
That is is correct

AZ:

The easiest way to translate words into an algebraic expression is to keep a few principles in mind: When you see the word 'is' that is the same thing as an equal sign If you see the word 'of' that is equivalent to multiplication The word 'sum' means addition The word 'difference' means subtraction For example, A number is the sum of oranges and apples can be translated to number = oranges + apples

AZ:

There's more phrases you can use to help you translate it into an algebraic expression which I will attach. And hopefully that will help you answer any future questions you have

1 attachment
mhanifa:

Do you still need help with this?